A compound is made up of A and B. A is placed on corner and face centre and B is placed on body centre and edge centre. What will be formula if all atoms on one body diagonal plane is removed???
How to do it??

Grade:12

1 Answers

Khimraj
5 years ago
initial no of atom
A = 8*1/8 + 6*1/2 = 4
B = 1*1 + 12*1/4 = 4
no of atom on one body diagonal
A = 2*1/8 = ¼
B = 1*1 = 1
remaining atom
A = 4 – ¼ = 15/4
B = 4 -1 = 4
so formula is A15/4B4 = A15B16
Hope it clears.